A stone is dropped from the top of tower touches the ground in 4 sec .the height of tower is about?

Grade:11

1 Answers

Satwik modi
6 years ago
As we know an object is dropped means it has initial velocity = 0
h=ut+1/2at^2
h=0*4+1/2 (10)*16
h=5*16
h=80 m